TopcoderArchiveVisit Topcoder
SRM26 Jul 2014
SRM 629

CandyDrawing

Problem statement, definition, constraints, and public examples.

Problem Statement

There are N boxes, numbered 1 through N. For each i, box i contains exactly i candies and one pebble. Thus, there are exactly i+1 objects in box i.

We are now going to create a collection of N objects using the following simple procedure: from each box, we will draw one object uniformly at random.

Let p be the probability that our collection will contain exactly K candies. You are given the ints N, K, and MOD. Return the value (p * (N+1)!) modulo MOD.

Definition

Class:
CandyDrawing
Method:
findProbability
Parameters:
int, int, int
Returns:
int
Method signature:
int findProbability(int N, int K, int MOD)
(be sure your method is public)

Constraints

  • N will be bewteen 1 and 1,000,000,000 (inclusive).
  • K will be between 0 and 2,000 (inclusive).
  • MOD will be between 1,000,000,000 and 2,000,000,000 (inclusive).
  • MOD will be a prime.

Examples

  1. 2
    1
    1000000007
    Returns: 3
    We have two boxes: box 1 with a candy and a pebble, and box 2 with two candies and a pebble. We are looking for the probability of drawing exactly one candy. This can happen in two different ways: either we draw the candy from box 1 and the pebble from box 2, or vice versa. The probability of the first way is (1/2)*(1/3) = 1/6. The probability of the second way is (1/2)*(2/3) = 1/3. Thus, the total probability is p = 1/6 + 1/3 = 1/2. We have p * (N+1)! = p*6 = 3, therefore the answer is (3 mod 1,000,000,007) = 3.
  2. 3
    2
    1000000007
    Returns: 11
  3. 10
    4
    1000000007
    Returns: 157773
  4. 1000000000
    1000
    1000000009
    Returns: 629516825
Back to all problems